OpenStudy (anonymous):

if an equation has an x and x^2 term how would you solve it give an example

OpenStudy (anonymous):

x^2-x=6 x^2-x-6=0 apply quadratic formula http://www.wolframalpha.com/input/?i=solve+x^2-x-6%3D0

OpenStudy (anonymous):

so the answer is apply the quadratic formula?

OpenStudy (anonymous):

yes sorry, i thought you had a question like cyter... which is right in that case (if you have an equation x²-x=6 then you can rearrange and apply the quadratic formula otherwise, if you have 4x * 3x² = 3 (which is what i thought of, then you could find the cube root) more likely your question is asking about what cyter said